From acd@roe.ac.uk Wed Jan 29 09:02:55 2014 Date: Tue, 28 Jan 2014 14:18:59 +0000 From: Clive Davenhall To: rgm@roe.ac.uk, al@roe.ac.uk, mar@roe.ac.uk, etws@roe.ac.uk, msh@roe.ac.uk, njc@roe.ac.uk, nch@roe.ac.uk, rpb@roe.ac.uk, rsc@roe.ac.uk, stv@roe.ac.uk, dmr@roe.ac.uk, acd@roe.ac.uk Cc: dante@astro.puc.cl, jsd@roe.ac.uk, j.p.emerson@qmul.ac.uk, keith@keithnoddle.org, Lorenzo.Rimoldini@unige.ch, mcioni@aip.de, matt.jarvis@astro.ox.ac.uk, mike@ast.cam.ac.uk, nigel.metcalfe@durham.ac.uk, norman@astro.gla.ac.uk, pwl@star.herts.ac.uk, rgm@ast.cam.ac.uk, Simon.Dye@nottingham.ac.uk, s.j.warren@ic.ac.uk, tkerr@jach.hawaii.edu, tom.shanks@durham.ac.uk Subject: WFAU VDFS Science Archives progress meeting minutes, 24 January 2014 WFAU VDFS Science Archives progress meeting minutes, 24 January 2014 Present: NJC, ETWS, MSH, RSC, RPB, ACD Apologies: MAR, RGM, NCH, KTN, AL DoNM: 10am Friday 7 February 2014 in the Villas meeting room ACD in the chair (MAR is away) (room is booked). Actions discharged this time: ---------------------------- None. Actions partly discharged but continuing: ---------------------------------------- ACTION: RSC to investigate modifying CU7 to recover ATLAS VST unmatched sources around frameset boundaries. Continues. Investigations had begun and a strategy had been worked out, but RSC was not sure how it would be implemented. He was in discussion with MAR about some of the details. Actions carried forward from 10/1/2014 meeting: ---------------------------------------------- ACTION: RPB to do the work necessary to make the pending VISTA public releases. Continues. VIKING has been done (DR3), VMC had fallen over and needed to be re-run. RPB had scheduled this work for yesterday but it was currently held up because of the problems with SVN. NEW ACTION: RPB, MSH investigate the of setting up a local nightly SVN backup. ACTION: ALL to add/maintain entries on http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Continues. ACD reiterated that he would be revising the entries shortly. ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Continues. ACTION: ACD to discuss with RGM the available budget for hardware purchases. Continues. Much progress had been made, but the amount was not yet finalised. ACTION: NJC to sort out the magnitude limits for V1.3 data. Continues. The software was largely written, but NJC had not had the time to finish and test it due to pressure of Gaia work. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. Specific points and new actions: ------------------------------- Project management: * Nothing to report. WFCAM, VISTA and VST updates: * Nothing to report. Comments and issues arising from CASU minutes: * No CASU minutes had been received since the last meeting. Networking: * ETWS reported: - Transfers to CASU of the hacked and deleted data are still continuing for WFCAM data. - Transfers from CASU of the latest released VISTA data for September and October 2013 are continuing. - The re-processed GPS files had been transferred (astrometry fix). WSA/VSA/OSA/GES Operations: * ETWS reported: - The VVV v1.3 data for P85, P86 and P87 are now fully ingested. The majority of P88 and P89 are also ingested, apart from some files where the ingests stopped due to various circumstances. Processing of P90 and P91 data is continuing. - Ingested most of the astrometry fixed GPS data, but found some that didn't ingest and sent them back to CASU. Mike Irwin will have a look when our transfer of the hacked WFCAM data back to them has finished. * No substantial work had been done on OSA or GES. Hardware and Systems: * MSH had upgraded Sahure and ETWS was installing the latest version of our 3rd-party software. Thoth would be done next week. * A purchase order had been placed for new backup software. * The SVN server was currently unavailable. * MSH needed to know an estimate of the budget available for hardware purchases as a matter of some urgency. ACTION: ACD to give MSH an estimate of the minimum amount available, as a matter of urgency. * Osorkon had been backed up and was ready to have its disks removed. Survey Data Release: * RPB noted: - He had largely prepared a VIKING P90 Consortium release, but still needed to do a few finishing touches. - Preparation of a VVV DR2 release was well-advanced and would be completed shortly. * NJC noted that he had been contacted by Omar Almaini about the date of the forthcoming UKIDSS DR11 release. RGM had asked AL to contact Steve Warren about the date of the release but a reply was still awaited. Omar would be visiting ROE next week and NJC would take this opportunity to discuss progress with him. Software: * NJC noted that he had moved some VMC-specific schema tables from one file to another. Previously some of the tables were in files that were not picked up by the release software and, after discussion with RSC, it was decided that it would be better to consolidate them into existing schema files where they would be picked up. * ETWS reported that he had fixed a small problem in which a file was still written-to while it was bulk ingested into the database. The fix comprised putting a wait into the CU1 code. Strangely this error happened only at around 2:45 during the last 2 nights, any other times the same code was executed, it worked without any problems. Non-survey Data Release: * There had not been any non-survey releases. * ETWS noted that recently no WFCAM data had been transferred from CASU. Astrogrid deployment, VO & Data Analysis services: * MSH had nothing to report, other than that DM and SV had visited last week. Miscellaneous: * ACD noted: - The RAS was soliciting for suggestions for Discussion Meetings. In the first instance anyone with any suggestions should try to convince RGM that they were a good idea. - The ASTROPLATE2014 workshop on the scientific use, digitisation and preservation of astronomical photographic records would be held in Prague on 18-21 March this year (note the short time-scale). ACD had details and anyone interested should contact him. * RSC was unable to attend the forthcoming ROE Jamboree. In the absence of any other volunteers ACD would present on his behalf. RSC then revealed that the presentation was to be given in traditional Cypriot nation dress. ACD, draft: 24/1/14, completed: 28/1/14. -- Scanned by iCritical.